4 along the y axix.
Perimeter=2l+2w
Consider the lines through P(2,4) and perpendicular to the x- and y-axes, respectively. Find the area and the perimeter of the rectangle formed by these lines and the axes.
I'm totally lost, please help. Thanks so much.

Answer by checkley77(12844)

The length of the rectangle is 2 along the x axix.
The height of the rectangle is 4 along the y axix.
Perimeter=2l+2w
Perimeter=2*2+2*4
Perimeter=4+8=12 ans.